Construct triangle EFG, if its perimeter is 9cm and angle F=60 degree and angle G=90 degree

Grade:7

1 Answers

sonika p
6 years ago
given GE+FG+FE=9cm....(1)
tan60=GE/FG = root 3 implies, FG=GE/root3
sin60= (root 3)/2=GE/FE implies FE=2GE/root3
putting all the values in equation 1, we get
GE=9/(root3 +1)
thus, FG=(3 * root3)/(1+root3)
FE=(6* root3)/(1+root3)